OpenStudy (dragon123mama):

i really hate geometry. it does not make as much sense to me as algebra did. please help. the midpoint of lineUV is (5,-11) The cordinates of one endpoint are u(3,5), find the cordinates of V?

OpenStudy (anonymous):

There are lots of ways to think about this. you can use the midpoint formula. \[(x,y)=\left( \frac{ x_1+x_2 }{ 2 },\frac{ y_1+y_2 }{ 2 } \right)\] Or you could think of it as 5 is halfway between 3 and what other number? -11 is halfway between 5 and what other number?

OpenStudy (dragon123mama):

i think personally formulas are my strong point., i think i just did not know what the formula was, thanks for reminding me

OpenStudy (anonymous):

you're welcome. You know how to set it up?

OpenStudy (dragon123mama):

yes once i see a formula and the numbers provided in the question it's easier for me to work it out.

OpenStudy (anonymous):

cool

OpenStudy (dragon123mama):

thanks